OpenStudy (anonymous):

what is the value of cosx if sinx=1/(1/square root of 3) in the range of 0

OpenStudy (anonymous):

is \[\sin x = \frac{ 1 }{ \sqrt{3} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use \[\sin ^{2}x + \cos ^{2}x = 1 \] to get the value of \[ cos x\]

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

\[\cos x = \sqrt{1-\sin ^{2}x}\] \[\cos x = \sqrt{1- \frac{ 1 }{ 3 }}\] \[\implies cosx = \sqrt{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

thankyou